# Off-Equatorial Orbits around Magnetically Charged Black Holes

### v1

@@ Abstract
We present a complete characterization of stable, off-equatorial circular orbits around magnetically charged black holes (MBHs). For a static, spherically symmetric MBH, we derive an exact analytic expression for the orbital latitude $\theta$ as a function of radius $r$. We establish a direct connection between these orbits and the spacetime fundamental structure, and demonstrate their stability against synchrotron radiation. We show that charged particles such as electrons and protons can exhibit $\mathcal{O}(1)$ latitude deviations at the ISCO radius and remain stable under synchrotron emission even for extremely small values of the black hole magnetic charge. We then extend the analysis to rotating MBHs, numerically computing the prograde and retrograde orbital branches and demonstrating how frame-dragging modifies their structure and stability regions. We show that these off-equatorial orbits are a unique feature of the magnetic charge, being forbidden in the analogous electrically charged Kerr-Newman spacetime. Our results suggest that environments surrounding magnetically charged black holes can exhibit distinctive phenomenological signatures, with potential implications for black hole imaging and polarimetric observations.
@@ I Introduction
The coupling of magnetic charge to gravity provides a rich theoretical laboratory for probing classical and semiclassical black-hole physics [MBH1, MBH2, Bai:2020spd, Bai:2020ezy, Gervalle:2024yxj]. The concept of magnetic monopole arises naturally in many extensions of the Standard Model, particularly in grand-unified theories (GUTs), where they are predicted to form as topological defects during symmetry-breaking phase transitions in the early universe [Dirac1931, tHooft1974, Polyakov1974]. Their gravitational counterparts, magnetically charged black holes (MBHs), serve as an idealized framework for studying how non-trivial gauge fields alter spacetime geometry and the dynamics within it, including the causal structure, particle motion, and radiative processes [Gibbons1975, Lee1992, Tursunov2018, Rayimbaev2021]. A key feature of these spacetimes is the non-zero divergence of the magnetic field, which prevents the electromagnetic vector potential from being defined globally. Nevertheless, a local potential can be defined in regions free of the monopole source, allowing for a consistent analysis of charged particle dynamics [Rohrlich1966, Preskill1984]. Furthermore, MBHs are connected to their electrically charged cousins—the Reissner–Nordström and Kerr–Newman black holes—via electromagnetic duality, a fundamental symmetry of Maxwell’s equations where the roles of electric and magnetic fields can be interchanged [duality1, duality2]. Studying the phenomenology of MBHs is therefore of intrinsic theoretical interest and may also inform searches for hypothetical observational signatures of such compact objects [Liu:2020vsy, Liu:2020bag, Chen:2022qvg] or distinguish them from exotic mimickers [Bah:2025vbr].
@@
The analysis of particle and light trajectories around black holes is a cornerstone of relativistic astrophysics [orbit1, orbit2, orbit3]. The paths of light are described by null geodesics, which determine the properties of gravitational lensing and define the location of the photon sphere—a region composed of unstable circular light orbits that is fundamental to the appearance of a black hole’s shadow [Synge1966, Bardeen1973, Chael2021]. The paths of massive objects are described by timelike geodesics, which govern the dynamics of accretion disks and define the innermost stable circular orbit (ISCO). The ISCO represents the minimum radius at which a particle can stably orbit a black hole before beginning an inevitable plunge, making it a critical parameter in models of accretion efficiency and compact-object inspirals [Bardeen1972, CompactInspiral, ISCO1, ISCO2]. The dynamics become significantly richer when the test particles are electrically charged. The Lorentz force, acting in concert with gravity, can produce qualitatively novel phenomena, such as strong latitudinal deflection and the emergence of stable, off-equatorial orbits not possible for neutral particles [Carter1968, Pugliese2011, Hackmann2013, Russo:2020lah, Dyson:2023ujk, Pereniguez:2024fkn]. A comprehensive analytic characterization of circular bound orbits around static and rotating MBHs, together with an analysis of their stability under electromagnetic emission and a connection between their properties and the fundamental structure of the spacetime, is still lacking in the literature. The present work aims to fill this gap.
@@
II Orbital Structure in Static MBH
@@
where the metric function $f(r)$ is given by Eq. (1) and $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\,d\phi^{2}$ is the line element on the unit 2-sphere. This metric is the magnetic analogue of the Reissner–Nordström spacetime and constitutes an exact solution to the Einstein–Maxwell equations for a purely radial magnetic field $\mathbf{B}=(P/r^{2})\,\hat{\mathbf{r}}$ [Reissner1916]. The cosmic censorship hypothesis imposes the bound $|P|\leq M$ [Penrose1965, Krolak1986], for which the spacetime possesses an event horizon ($r_{+}$) and an inner Cauchy horizon ($r_{-}$) at
@@
Due to the magnetic charge, the electromagnetic four-potential $A_{\mu}$ cannot be globally defined. It can, however, be expressed locally on gauge patches. For the northern hemisphere ($0\leq\theta<\pi$), we adopt the standard form [Rohrlich1966, Preskill1984]:
@@
The motion of a test particle with rest mass $m$ and electric charge $q$ is governed by the the geodesic equation, for a neutral particle ($q=0$) it gives:
$\frac{d^{2}x^{\mu}}{d\tau^{2}}+\Gamma^{\mu}_{\alpha\beta}u^{\alpha}u^{\beta}=0.$ (5)
For a charged particle, the equation of motion includes the Lorentz force:
where $\Gamma^{\mu}_{\alpha\beta}$ are the Christoffel symbols.
For neutral particles, the spacetime symmetries guarantee the conservation of energy $E$ and angular momentum $J$. For photons ($m=0,q=0$), which satisfy the null condition $g_{\mu\nu}u^{\mu}u^{\nu}=0$, circular orbits define the photon sphere. Its radius, $r_{\text{ps}}$, is found by extremizing the effective potential, yielding
@@
which yields the cubic equation
@@
The presence of electric charge of the test particle introduces a Lorentz force that alters the orbital plane, displacing the trajectory from the equator. Stable circular orbits still exist for certain polar angles, as shown in Figure 1.
@@
For a test particle moving on a circular orbit, its 4-velocity is given by
$u^{\mu}=(u^{t},0,0,u^{\varphi}),$ (11)
with the constraint
$u^{\mu}u_{\mu}=-f(r)(u^{t})^{2}+\frac{\dot{r}^{2}}{f(r)}-r^{2}(u^{\varphi})^{2}=-1.\\$ (12)
The $t$ and $\varphi$ components of Eq. (6) vanish identically as
@@
$\frac{d\varphi}{d\tau}=\omega_{\pm}=\pm\frac{1}{r\sin\theta}\sqrt{\frac{\frac{M}{r}-\frac{P^{2}}{r^{2}}}{1-\frac{3M}{r}+\frac{2P^{2}}{r^{2}}}}.$ (15)
@@
For the $\varphi$ equation, the electromagnetic contribution yields
$F^{\mu\nu}u_{\nu}=g^{\mu\alpha}g^{\nu\beta}F_{\alpha\beta}u_{\nu}=\frac{P}{r^{2}}\omega\sin\theta,$
@@
$\Gamma^{\varphi}_{\alpha\beta}\frac{dx^{\alpha}}{d\tau}\frac{dx^{\beta}}{d\tau}=\Gamma^{\varphi}_{\varphi\varphi}u^{\varphi}u^{\varphi}=-\omega\sin\theta\cos\theta.$
@@
This expression indicates that a realistic particle, such as an electron with a charge-to-mass ratio of $\sim 10^{22}$, can still have $\tan(\theta)\sim\mathcal{O}(1)$ for values of $P/M$ as small as $10^{-22}$, highlighting the strong influence that even a tiny magnetic charge can exert on light charged particle. This is somehow reminiscent of the findings in Ref. [Pereniguez:2024fkn], where the authors show that even a single magnetic monopole in the black hole can significantly affect the superradiant instability timescale.
@@
Figure 2: $\theta_{+}(r)$ of a q/m=1q/m=-1 particle, for multiple $P/M$ MBHs. $\theta_{+}<\pi/2$ since the magnetic force on a prograde, negatively charged particle points above the equator. The y-axis is inverted so that north pole direction ($\theta=0$) appears at the top of the figure, matching the intuitive orientation.
@@
Figure 3 shows the upper half of a cross section of a MBH with $P/M=0.6$, together with representative prograde circular orbits for test particles with q/m=1q/m=-1 and q/m=10q/m=-10. The curves in the $(r,\theta)$ plane indicate a continuous series of such orbits. The black disk marks the horizon, while the grey and blue dashed circles denote the photon sphere and the ISCO radius, respectively. This is a striking consequence of the long-range magnetic field.
@@
III Analysis of Orbits in the Static Spacetime
We now perform a detailed analysis of the off-equatorial circular orbits derived in the previous section, connecting their properties to the fundamental features of the spacetime and assessing their stability.
@@ III Analysis of Orbits in the Static Spacetime
Next, we turn to the innermost stable circular orbit (ISCO) of a charged particle. Unlike the neutral case, the Lorentz force alters the direction of the tangential velocity, so the orbit cannot remain confined to the equatorial plane. Thus, the 4-velocity must be written as
$u^{\mu}=(\dot{t},\dot{r},\dot{\theta},\dot{\varphi}),$
and the normalization condition becomes
The corresponding particle Lagrangian is
$\mathcal{L}=-\frac{m}{2}u^{\mu}u_{\mu}+qu^{\mu}A_{\mu}=-\frac{m}{2}u^{\mu}u_{\mu}+qA_{\varphi}\dot{\varphi}.$ (21)
Time independence of the Lagrangian guarantees conservation of energy:
@@
III.1 Synchrotron radiation of charged particles
Finally, we analyze the effect of synchrotron radiation on circular orbits, assessing the conditions under which these orbits can persist for sufficiently long lifetimes. In curved spacetime, the equation of motion of a charged particle incorporating the radiation-reaction effects reads [Bracken2024]
@@ III.1 Synchrotron radiation of charged particles
The existence of a timelike Killing vector $k^{\mu}=(1,0,0,0)$ allows us to define the energy of the particle as $E=-m\,k_{\mu}\dot{z}^{\mu}$. The emitted power is then found by computing
@@
and the definitions of $\sigma$, $u_{\mu\nu}$ $w_{\mu\nu}$ can be found is Refs. [Najmi1985, DeWitt1960]. In the coincidence limit one has [DeWitt1960]:
@@
indicating that Eq. (36) yields a negligible contribution. Analogously, one can verify that $(\dddot{z})_{0}=0$ along circular obits. Therefore, the only non-vanishing terms in Eq. (35) are
@@
To visualize the parameter space where orbits are both stable and non-trivial, we plot the constraints in Figure 4. The dashed lines indicates the regime where the magnetic field is strong enough to cause a significant latitudinal deviation ($|\tan\theta|<1$), as expressed in Eq. (17). The red region delineates the area of parameter space where the ratio $\delta\tau_{\text{rad}}/t_{\text{orb}}<1$, for a benchmark value of $m$ equal to the electron mass and $M=10^{6}\,M_{\odot}$. For smaller values of $m/M$, the dashed red line shifts even upward in Figure 4, whereas the opposite trend occurs if we increase $m/M$.
Figure 4: The parameter space of circular orbits for a charged particle with charge-to-mass ratio $|q/m|$ orbiting a massive black hole (MBH) with charge-to-mass ratio $|P/M|$. The dashed lines indicates where the circular orbit exhibits a significant latitudinal deviation ($|\tan\theta|<1$). The red region marks the values of the parameters for which the orbit becomes unstable under the emission of synchrotron radiation. The extent of this unstable region depends on the ratio $m/M$, as shown in Eq. (46). In this figure, we take $m$ to be the electron mass and $M=10^{6}\,M_{\odot}$.
IV Orbital structure in Rotating MBH
The charged spinning spacetime is described by the Kerr–Newman metric [Kerr1963, Mazur1982], which is given by

### v2

@@ Abstract
We present a complete characterization of stable, off-equatorial circular orbits around magnetically charged black holes (MBHs). For a static, spherically symmetric MBH, we derive an exact analytic expression for the orbital latitude $\theta$ as a function of radius $r$ and we analyze the effect of synchrotron radiation. We show that charged particles such as electrons and protons can exhibit $\mathcal{O}(1)$ latitude deviations at the ISCO radius and remain stable under synchrotron emission even for extremely small values of the black hole magnetic charge. We then extend the analysis to rotating MBHs, numerically computing the prograde and retrograde orbital branches and demonstrating how frame-dragging modifies their structure and stability regions. We show that these off-equatorial orbits are a unique feature of the magnetic charge, being forbidden in the analogous electrically charged Kerr-Newman spacetime. Our results suggest that environments surrounding magnetically charged black holes can exhibit distinctive phenomenological signatures, potentially offering a way to constrain the magnetic charge of astrophysical black holes.
@@ I Introduction
The coupling of magnetic charge to gravity provides a rich theoretical laboratory for probing classical and semiclassical black-hole physics [29, 33, 4, 5, 21]. Magnetic monopole arise naturally in many extensions of the Standard Model, particularly in grand-unified theories (GUTs), where they are predicted to form as topological defects during symmetry-breaking phase transitions in the early universe [15, 1, 40]. Their gravitational counterparts, magnetically charged black holes (MBHs), serve as an idealized framework for studying how non-trivial gauge fields alter spacetime geometry and the dynamics within it, including the causal structure, particle motion, and radiative processes [22, 30, 51, 43]. A key feature of these spacetimes is the non-zero divergence of the magnetic field, which prevents the electromagnetic vector potential from being defined globally. Nevertheless, a local potential can be defined in regions free of the monopole source, allowing for a consistent analysis of charged particle dynamics [45, 41]. Furthermore, MBHs are connected to their electrically charged cousins—the Reissner–Nordström and Kerr–Newman black holes—via electromagnetic duality, a fundamental symmetry of Maxwell’s equations where the roles of electric and magnetic fields can be interchanged [25, 39]. Studying the phenomenology of MBHs is therefore of intrinsic theoretical interest and may also inform searches for observational signatures of such compact objects [31, 32, 13, 23] or distinguish them from exotic mimickers [3].
@@
The analysis of particle and light trajectories around black holes is a cornerstone of relativistic astrophysics [18, 48, 20]. The paths of light are described by null geodesics, which determine the properties of gravitational lensing and define the location of the photon sphere, a region composed of unstable circular light orbits that is fundamental to the appearance of a black hole’s shadow [50, 6, 12]. The paths of massive objects are described by timelike geodesics, which govern the dynamics of accretion disks and define the innermost stable circular orbit (ISCO). The ISCO represents the minimum radius at which a massive particle can stably orbit a black hole before beginning an inevitable plunge, making it a critical parameter in models of accretion efficiency and compact-object inspirals [7, 47, 49, 52]. The dynamics become significantly richer when the test particles are electrically charged. The Lorentz force, acting in concert with gravity, can produce qualitatively novel phenomena, such as strong latitudinal deflection and the emergence of stable, off-equatorial orbits not possible for neutral particles [11, 42, 24, 46, 16, 38]. A comprehensive analytic characterization of circular bound orbits around static and rotating MBHs, together with an analysis of their stability under electromagnetic emission, is still lacking in the literature. The present work aims to fill this gap.
@@
II Orbital Structure in Static MBHs
@@
where the metric function $f(r)$ is given by Eq. (1) and $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\,d\phi^{2}$ is the line element on the unit 2-sphere. This metric is the magnetic analogue of the Reissner–Nordström metric and constitutes an exact solution to the Einstein–Maxwell equations for a purely radial magnetic field $\mathbf{B}=(P/r^{2})\,\hat{\mathbf{r}}$ [44]. The cosmic censorship hypothesis imposes the bound $|P|\leq M$ [37, 28], for which the spacetime possesses an event horizon ($r_{+}$) and an inner Cauchy horizon ($r_{-}$) at
@@
Due to the magnetic charge, the electromagnetic four-potential $A_{\mu}$ cannot be globally defined. It can, however, be expressed locally on gauge patches. For the northern hemisphere ($0\leq\theta<\pi$), we adopt the standard form [45, 41]:
@@
The spacetime possesses two Killing vectors: a timelike Killing vector $k^{\mu}=(1,0,0,0)$ associated with time-translation invariance, and an azimuthal Killing vector $\xi^{\mu}=(0,0,0,1)$ associated with rotational symmetry. For neutral particles, these symmetries guarantee the conservation of energy $E$ and angular momentum $J$.
For photons ($m=0,q=0$), which satisfy the null condition $g_{\mu\nu}u^{\mu}u^{\nu}=0$, circular orbits define the photon sphere. Its radius, $r_{\text{ps}}$, is found by extremizing the effective potential, yielding
@@
which, after solving for $j^{2}$, yields the cubic equation
@@
For a test particle with rest mass $m$ and charge $q$, its 4-velocity
$u^{\mu}=(\dot{t},\dot{r},\dot{\theta},\dot{\varphi}),$ (9)
satisfies the constraint
Where the over-dot denotes a derivative with respect to proper time $\tau$. The equations of motion can be derived from the Lagrangian
$\mathcal{L}=-\frac{m}{2}u^{\mu}u_{\mu}+qu^{\mu}A_{\mu}=-\frac{m}{2}u^{\mu}u_{\mu}+qA_{\varphi}\dot{\varphi},$ (11)
which leads to
where $\Gamma^{\mu}_{\alpha\beta}$ are the Christoffel symbols.
The presence of electric charge of the test particle introduces the Lorentz force that alters the orbital plane, displacing the trajectory from the equator. Despite this, stable circular orbits still exist for certain polar angles, as shown in Figure 1.
For the test particle to move on a circular orbit, its 4- velocity simplifies to
$u^{\mu}=(\dot{t},0,0,\dot{\varphi}),$ (13)
and $f(r)\dot{t}^{2}-r^{2}\dot{\varphi}^{2}=1$. The $t$ and $\varphi$ components of Eq. (12) vanish identically as
@@
$u^{\varphi}=\omega_{\pm}=\pm\frac{1}{r\sin\theta}\sqrt{\frac{\frac{M}{r}-\frac{P^{2}}{r^{2}}}{1-\frac{3M}{r}+\frac{2P^{2}}{r^{2}}}}.$ (16)
@@
For the $\theta$ equation, the electromagnetic contribution yields
$\frac{q}{m}F^{\theta\nu}u_{\nu}=\frac{q}{m}g^{\theta\alpha}g^{\nu\beta}F_{\alpha\beta}u_{\nu}=\frac{q}{m}\frac{P}{r^{2}}\omega\sin\theta,$
@@
$\Gamma^{\theta}_{\alpha\beta}\frac{dx^{\alpha}}{d\tau}\frac{dx^{\beta}}{d\tau}=\Gamma^{\theta}_{\varphi\varphi}u^{\varphi}u^{\varphi}=-\omega^{2}\sin\theta\cos\theta.$
@@
This expression indicates that a realistic particle, such as an electron with a charge-to-mass ratio of $\sim 10^{22}$, can still have $\tan(\theta)\sim\mathcal{O}(1)$ for values of $P/M$ as small as $10^{-22}$, highlighting the strong influence that even a tiny magnetic charge can exert on light charged particle. This is somehow reminiscent of the findings in Ref. [38], where the authors show that even a single magnetic monopole in the black hole can significantly affect the superradiant instability timescale.
@@
Figure 2: $\theta_{+}(r)$ of a q/m=1q/m=-1 particle, for multiple $P/M$ MBHs. Solid lines denote stable orbits, while dashed lines indicate instability against small perturbations; the transition point (ISCO) is marked on each curve. Note that $\theta_{+}<\pi/2$ since the magnetic force on a prograde, negatively charged particle points above the equator. The y-axis is inverted so that north pole direction ($\theta=0$) appears at the top of the figure, matching the intuitive orientation.
@@
Figure 3 shows the upper half of a cross section of a MBH with $P/M=0.6$, together with representative prograde circular orbits for test particles with q/m=1q/m=-1 and q/m=10q/m=-10. The curves in the $(r,\theta)$ plane indicate a continuous series of such orbits. The black disk marks the horizon, while the grey and blue dashed circles denote the photon sphere and the ISCO radius, respectively.
@@
III Analysis of Orbits in Non-Rotating MBHs
We now perform a detailed analysis of the off-equatorial circular orbits derived in the previous section and assess their stability.
III.1 Radial Boundaries and Stability Analysis
@@
Next, we investigate the innermost stable circular orbit (ISCO) of a charged particle. The time independence of the Lagrangian (Eq. (11)) guarantees conservation of energy:
@@
III.2 Synchrotron radiation of charged particles
Finally, we analyze the effect of synchrotron radiation on circular orbits, assessing the conditions under which these orbits can persist for sufficiently long lifetimes. In curved spacetime, the equation of motion of a charged particle incorporating the radiation-reaction effects reads [9]
@@ III.2 Synchrotron radiation of charged particles
The timelike Killing vector $k^{\mu}=(1,0,0,0)$ introduced in Section II allows us to define the energy of the particle as $E=-m\,k_{\mu}\dot{z}^{\mu}$. The emitted power is then found by computing
@@
and the definitions of $\sigma$, $u_{\mu\nu}$ $w_{\mu\nu}$ can be found is Refs. [35, 14]. In the coincidence limit one has [14]:
@@
indicating that Eq. (35) yields a negligible contribution.
Analogously, one can verify that $(\dddot{z})_{0}=0$ along circular obits. Therefore, the only non-vanishing terms in Eq. (34) are
@@
Figure 4: The parameter space of circular orbits for a charged particle with charge-to-mass ratio $|q/m|$ orbiting a massive black hole (MBH) with charge-to-mass ratio $|P/M|$. The dashed yellow lines indicate where the circular orbit exhibits significant latitudinal deviations ($|\tan\theta|=1,10^{-3},10^{-6}$). The red region marks the values of the parameters for which the orbit becomes unstable under the emission of synchrotron radiation. The extent of this unstable region depends on the ratio $m/M$, as shown in Eq. (45). In this figure, we take $m$ to be the electron mass and $M=10^{6}\,M_{\odot}$.
@@
To visualize the parameter space where orbits are both stable and non-trivial, we plot the constraints in Figure 4. The dashed yellow lines correspond to constant values of $|\tan\theta|$, quantifying the latitudinal deviation caused by the magnetic field as per Eq. (18). Specifically, smaller values of $|\tan\theta|$ indicate orbits situated further from the equatorial plane. The red region delineates the area of parameter space where the timescale ratio $\delta\tau_{\text{rad}}/t_{\text{orb}}<1$, marking the system as unstable to synchrotron radiation. This boundary is calculated for a benchmark system with $m$ equal to the electron mass and $M=10^{6}\,M_{\odot}$. Decreasing the mass ratio $m/M$ shifts the instability boundary (solid red line) upward in Figure 4, whereas the opposite trend occurs if we increase $m/M$.
IV Orbital structure of Rotating Black Holes
In this section, we extend our analysis to rotating black holes. The angular momentum of the black hole breaks spherical symmetry, reducing it to axial symmetry. Thus, we naturally align the $z$-axis with its rotation axis.
IV.1 Circular Orbits in Rotating MBHs
The charged spinning spacetime is described by the Kerr–Newman metric [27, 34], which is given by